OpenStudy (luigi0210):

Integration:

OpenStudy (luigi0210):

\[\LARGE \int~\frac{dx}{5+4cosx}\]

OpenStudy (inkyvoyd):

can I just call this a poopy problem and not try to do it?

geerky42 (geerky42):

Who told you to do it? lol

OpenStudy (anonymous):

put \[\tan \frac{ x }{ 2 }=t,x=2~\tan^{-1} t,dx=\frac{ 2 }{ 1+t^2 }dt\] \[\cos x=\frac{ 1-\tan ^2\frac{ x }{ 2 } }{ 1+\tan ^2\frac{ x }{ 2 } }\]